f(x)=(sq(m)-1))xsq + (m-1)x+n+2(m not equal to 1)


be an odd function and m and n are constants.determine whether g(x)=x(power)m +x (POWer)n is an even or odd or none.

Grade:12

2 Answers

manoj jangra
14 years ago

hey its answer is none .f(x) = (m- 1)x2 + (m - 1)x + n + 2     

for odd function f(x) = - f(x)................................................(1)

here as xis even function its coffeceiont must be 0

m- 1 = 0

m = -1   

  after solving equation 1 we get n=-2

g(x) will be neither even nor odd   

 

abhishek upadhyay
12 years ago

dear chauhan

  quadratics are never odd they are even if coeff. of x is 0